a woman cycles for x hours at 12km/h and y hours at 18km/h. altogether she cycles 78km in 5 hours. find x and y?
Answer by mananth(16946)

Total time
x+y=5
Total distance
12x+18y=78
1.00 x + 1.00 y = 5.00 .............1
Total value
12.00 x + 18.00 y = 78.00 .............2
Eliminate y
multiply (1)by -18.00
Multiply (2) by 1.00
-18.00 x -18.00 y = -90.00
12.00 x + 18.00 y = 78.00
Add the two equations
-6.00 x = -12.00
/ -6.00
x = 2.00
plug value of x in (1)
1.00 x + 1.00 y = 5.00
2.00 + y = 5.00
y = 5.00 -2.00
y = 3.00
y = 3.00
x= 2.00
y= 3.00
m.ananth@hotmail.ca